North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Home Locomotives Q6 63395 (NER T2 2238) Q6 Blog Saturday 24th March 2018

Saturday 24th March 2018

Ian reports a grey wet morning at Grosmont yesterday, but the sun did shine after lunch. Peter Ellis, Alan Hardie, Bill Dobson, and Ian were present for the usual cuppa, and Andy Lowes joined after his PTS course finished.

Work started with Peter making and fitting cotters to the eccentrics and Ian fitting taper pins to the eccentric rods. Bill fitted the R/H trailing wheel crank pin boss after reaming it first, then fitting the side rod end caps and taper pins: only the driving wheels to do when the L/H big end brass comes back from the machine shop, promised for Wednesday. Alan was adjusting the large spring hanger nuts, so just the driving springs left to do at the moment. Peter completed fitting cotters and Ian completed fitting the taper pins.

Lunch was had upstairs at the MPD where Andy Lowes and Steve Hyman joined the team after their PTS course. After lunch Andy came up to Deviation Shed and did a series of checks on the work that had been done on taper pins, cotters, split pins and various fittings. He pointed out that there were some split pins to open and nuts to tighten: these have been noted for attention on Wednesday. Peter started fitting the pressure gauges in the cab. The steam pressure and steam heat gauges were fitted but not connected to their copper pipes. The vacuum gauge was taken back to the MPD office as there were no screws holding the outer case on: it was left on Barney's desk with a note from Bill regarding the defect. The compressor was also turned on and the air line connected to the reverser to try and move the valve rod back sufficiently to remove the L/H follower, but, after several attempts, it would not move either way, even with it well oiled and the gland nuts slackened off. The hope is that we don't have to strip it down!

Two flue tube have been given to us for our pot bellied stove by Kevin Yeoman of the Dame Vera Lynn Group, and are outside the east side of Deviation Shed awaiting fitting to the stove. Our thanks to Kevin for these.

Quite a few visitors yesterday now the trains are running again, or at least trying to! The B1 on the first train struggled past Deviation Shed and only got past Esk Valley Cottages where it was failed with a faulty pressure gauge which was giving a false reading. It was rescued by No 926 Repton which then slipped and slid its way to Goathland with the next train, causing more late running. Things did get better later on though, with the B1 repaired and back in traffic to haul the 16.30 last train to Pickering.

Working party next Wednesday as usual. Red timetable in operation. Hope to see you there.